A car manufacturer claims that the variance of the gas mileage in a model of it new release self-driving car is more than 0.24. You work for consumer advocacy group and are responsible for testing this claim. you sampled 23 vehicles and came up with a variance of 0.17 instead. At alpha = 0.1, is there enough evidence to reject the claim? Assume test scores are normally distributed.
Answer #1

Ans:

Ho : σ2 < 0.24

H, : 0 > 0.24

Test statistic:

Chi square=(23-1)*0.17/0.24=15.583

df=23-1=22

p-value=CHIDIST(15.583,22)=0.8360

As,p-value>0.1,we fail to reject the null hypothesis.

There is not sufficient evidence to support the claim that the variance of the gas mileage in a model of it new release self-driving car is more than 0.24